OpenStudy (anonymous):

integral by parts: ʃ -cosx e^9x

OpenStudy (amistre64):

this should circle back on itself

OpenStudy (anonymous):

?

OpenStudy (amistre64):

ill see your ? and raise you a !!

OpenStudy (anonymous):

sorry?

OpenStudy (amistre64):

integrate this by parts twice and you should be able to circle back on yourself to get rid of the cycle

OpenStudy (amistre64):

unless of course you havent even tried it yet, and thereby have no idea what the issue is

OpenStudy (anonymous):

i got -1/9 e^9x cosx-1/9e^9x cosx

OpenStudy (anonymous):

is that correct?

OpenStudy (amistre64):

dunno, lets see if i can come up with something that doesnt bog down my browser to a crawl :)

OpenStudy (turingtest):

can't be right they are both cos...

OpenStudy (amistre64):

{S} -cosx e^9x dx = -cos(x) + e^9x sin(x) - 9 e^9x cos(x) + 81 e^9x ...... {S} -cosx e^9x dx = e^9x sin(x) -9 e^9x cos(x) - {S} 81 cos(x) e^9x dx -{S} cosx e^9x dx = e^9x sin(x) -9 e^9x cos(x) - 81{S} cos(x) e^9x dx + 81{S} cos(x) e^9x dx + 81{S} cos(x) e^9x dx ---------------------------------------------------------------- 80 {S} cosx e^9x dx = e^9x sin(x) -9 e^9x cos(x) /80 /80 ---------------------------------- {S} cosx e^9x dx = (e^9x sin(x) -9 e^9x cos(x))/80 maybe lol

OpenStudy (amistre64):

close, real close lol

OpenStudy (anonymous):

lol not convinced

OpenStudy (amistre64):

i got some sign issues, but other than that :)

OpenStudy (amistre64):

i see it, I was deriving my -cos instead of integrating it; so my signs got off lol

OpenStudy (anonymous):

confusing :)

OpenStudy (amistre64):

yeah, but the whole issue is in that it tends to repeat itself with integration by parts so you have to break that cycle

OpenStudy (anonymous):

yh thanks..what grade r u in

OpenStudy (amistre64):

ran out of grades; 35th maybe? i lost count

OpenStudy (anonymous):

what do u mean?

OpenStudy (amistre64):

it means that at my age, you dont have a grade to be in .....

OpenStudy (anonymous):

oh ok u must be way past university level??

OpenStudy (amistre64):

must be ;)

OpenStudy (anonymous):

Dr professor?

OpenStudy (anonymous):

usa?